# Talking Heads: Understanding Inter-layer Communication in Transformer Language Models

### v1

@@ Abstract
Although it is known that transformer language models (LMs) pass features from early layers to later layers, it is not well understood how this information is represented and routed by the model. By analyzing particular mechanism LMs use to accomplish this, we find that it is also used to recall items from a list, and show that this mechanism can explain an otherwise arbitrary-seeming sensitivity of the model to the order of items in the prompt. Specifically, we find that models write into low-rank subspaces of the residual stream to represent features which are then read out by specific later layers, forming low-rank communication channels between layers. By decomposing attention head weight matrices with the Singular Value Decomposition (SVD), we find that previously described interactions between heads separated by one or more layers can be predicted via analysis of their weight matrices. We show that it is possible to manipulate the internal model representations as well as edit model weights based on the mechanism we discover in order to significantly improve performance on our synthetic Laundry List task, which requires recall from a list, often improving task accuracy by over 20%. Our analysis reveals a surprisingly intricate interpretable structure learned from language model pretraining, and helps us understand why sophisticated LMs sometimes fail in simple domains, facilitating future analysis of more complex behaviors.
@@ 1 Introduction
Despite the impressive capabilities of LMs, their practical use is often limited because of seemingly arbitrary sensitivities to prompts. These failure cases are particularly troubling because they are not systematic; it is very difficult to predict when, for example, the order of information seemingly randomly causes a model to fail [Pezeshkpour and Hruschka, 2023, Liu et al., 2024, Li and Gao, 2024, Zheng et al., 2024, Zhou et al., 2023], or the format of a prompt hurts performance [Liu et al., 2023, Sclar et al., 2023, Zhao et al., 2021, Lu et al., 2022, Webson and Pavlick, 2022]. As LLMs become increasingly ubiquitous, we will require more principled ways of anticipating and remedying unstable or unwanted behaviors [Yu et al., 2024, Yong et al., 2023]. Understanding the mechanisms in play within LLMs, and connecting those mechanisms to behavior, could enable such principled approaches.
@@
In this paper, we consider a simple laundry list task that exhibits one such undesirable instability. Specifically: Transformer language models (LMs) struggle to reliably recall items from a list as the length of the list increases, and performance can vary wildly depending on the position of the item in the list that is being recalled (Figure 1). This instability is not obvious from the model architecture itself–i.e., unlike their predecessors Elman [1991], Transformers [Vaswani et al., 2017] do not have built-in biases for some positions in a sequence over others. We thus use this task as a case study in order to connect the low-level emergent mechanisms which are encoded during LM pretraining to observable behavior, and illustrate as a proof of concept that a precise mechanistic understanding of LMs can be used to explicate and, perhaps, remedy model performance in practice.
@@
Specifically, building on recent work in circuit analysis [Elhage et al., 2021, Wang et al., 2022, Goldowsky-Dill et al., 2023, Quirke and Barez, 2024, Merullo et al., 2024, Hanna et al., 2023], we demonstrate how a Transformer LM (GPT2 small) passes information from early laters to later ones using subspaces which we call communication channels. We introduce a method using the Singular Value Decomposition to help find these channels by reading off of the model weights directly. We focus on two examples of such channels (inhibition and duplicate detection), and find that they are are very low rank (1 or 2 dimensions), easily interpretable, and causally important for specific model behaviors. Specifically, our contributions are as follows:
@@
  • We explore a fundamental question in interpretability on how information passed from layer to layer in a transformer is represented internally. We find “communication channels" that connect attention heads separated by several layers.
  • We introduce a method to decompose weight matrices which helps us find such communication channels. This method can be performed directly on the weights of models.
@@
  • We show that this mechanism plays a role in prompt sensitivity on an item recall task that otherwise seems idiosyncratic, and that intervention in the mechanism can be used to affect downstream behavior in a way that improves performance.
@@
Our findings indicate more broadly that neural networks are capable of learning intricately structured representations from self-supervised pretraining without inductive biases, which might have important implications for the emergence of abstract and content-independent operations, and for developing methods for steering and understanding these models (see Discussion, §7).
@@
Figure 1: Language models are sensitive to arbitrary changes in a prompt, for example the order in which objects are listed (right). This problem is more pronounced as the number of objects increases (left) even though it is not obvious where the issue stems from in the model. We broadly explore how information is routed through a model and focus on a mechanism that is in part responsible for this (in)ability.
@@ 2 Background
Throughout our work, we consider decoder-only transformer language models and primarily use GPT2-Small. The modern attention mechanism used by these models Vaswani et al. [2017] uses multiheaded attention, where Query and Keys control which tokens from earlier in context are attended to and Value and Output matrices control what information is moved from these tokens. An important abstraction we use in our work relies on rewriting these matrices as the products QK (Query*Key) and OV (Output*Value) [^1] Following convention, we refer to this matrix as OV, but the Transformer Lens library implements right-hand matrix multiplication so we actually use VO. This does not effect our results. For an individual head, these matrices are themselves low-rank compared to the embedding dimension of the network. This is a useful property as we are motivated by looking for subspaces that are written into/read from by these matrices and this reduces the search space.
@@
We build on work from IOI [Wang et al., 2022] which documents a circuit that appears in multiple tasks [Merullo et al., 2024]. This circuit includes mover heads, which copy tokens from context to the output, and inhibition heads which (optionally) block the mover heads’ attention to certain tokens and thus prevent certain tokens from being copied. Inhibition heads are known to receive signals from duplicate token head value vectors which help inform which tokens to inhibit. In GPT2-Small, the known inhibition heads are 7.3, 7.9, 8,6, and 8.10 and we consider their communication to the mover head 9.9222for simplicity we only consider this mover head, but we do not find the choice matters much.. This is the example we use for query composition experiments. In Section 5 we explore this circuit’s role in problems of prompt sensitivity and a learned structure to control the indexing of tokens in the context window.
@@ 3 Identifying Communication Channels in Low-Rank Subspaces
In this section, we test the hypothesis that model components like attention heads communicate through signals in low-rank subspaces of the residual stream and that we can find these signals in the weights themselves. We investigate one case of each type of composition outlined in Section 2 and find positive evidence for this hypothesis with query and value composition in 1 and 2 dimensional subspaces, but not key composition (see Appendix A). Because we are able to localize the query and value signals to such small representation spaces, we find that we are able to interpret and control these features with intervention experiments in Section 4.
@@ 3.1 Composition Score
The Composition Score (CS) is a weight based metric of how much two weight matrices ‘talk’ to each other when they are separated by layers. That is, $\mathbf{W_{1}}$ might write information into a subspace in the residual stream that is read out by $\mathbf{W_{2}}$. In Query composition for example, $\mathbf{W_{1}}$ is the OV matrix of some head, and $\mathbf{W_{2}}$ is the QK matrix.
@@
We take advantage of the fact that circuit analysis in works like Wang et al. [2022] tells us that, for example, head 3.0 (duplicate head) interacts with head 7.9 (inhibition head) through value composition and 7.9 with 9.9 (mover head) in query composition. We initially use the composition score in Equation 1 to try predict these interactions in the weights, but find these results are largely noisy and uninterpretable. This is briefly demonstrated in Appendix G. We find that despite empirically knowing interactions exist between heads, we do not find they reliably have higher composition scores than any random head. Although the composition score has been shown to be useful on small toy models [Elhage et al., 2021], previous work has also shown that on larger models, the signal the composition score conveys is extremely noisy [Singh et al., 2024].
@@ 3.2 Composition with Decomposed Component Matrices
The composition is not useful on its own because otherwise important signals are possibly washed out by the fact that every head reads from/writes into many subspaces to some extent. Therefore, we turn to the Singular Value Decomposition (SVD), defined as $\mathbf{W}=\mathbf{USV^{T}}$, on the QK and OV matrices to decompose the attention heads into orthogonal components which determine the input and output spaces of the matrix. This allows us to individually view subspaces read from/written into ordered by the amount of variance of the transformation of the matrix they account for. This helps us answer our original hypothesis that model components communicate across layers in low-rank subspaces of the residual stream.
@@
We find that decomposing weight matrices this way is very effective at cleaning up the composition score signal. We find attention heads that have very high relative composition scores with one component matrix of another head. For example, the second component of head 8.6 (referred to as 8.6.2) composes far more with mover head 9.9 than any of the other 63 components in 8.6 (5 standard deviations higher than the average) or when considering the full matrix as in $CS(\mathbf{OV}^{8.6},\mathbf{QK}^{9.9})$. The bottom graphs in Figure 2 show these results for the inhibition heads. All of the inhibition head exhibit a similar phenomenon of single component dominance. The duplicate token head 3.0 also value-composes with inhibition heads in a similar way, using two components (3.0.1 and 3.0.2) far more than any other. Results are shown in Appendix C.
We can also use this decomposition to find specific pairs of heads that talk more than others. With the knowledge that two heads talk through a specific component of one head, we can find the other heads that communicate through this pathway. Doing so lets us find the signal encoding almost the full IOI circuit in GPT2-Small directly from the weights, without running the model. We outline these results in Appendix G.
@@ 3.3.2 Results
Our editing technique is simply to zero out one component at a time and across a dataset of IOI examples test how this affects. One way to think about this is zeroing out one singular value of e.g., the OV or matrix, or subtracting one of the component matrices from the sum in Equation 2. We must make the edit to the decomposition and then split the matrices back out so that we can run the model. Given $\mathbf{OV}=\mathbf{USV}^{T}$, after zeroing out some singular value in $\mathbf{S}$ we can set the Output and Value matrices to be $\mathbf{U}\sqrt{\mathbf{S}}$ and $\sqrt{\mathbf{S}}\mathbf{V}^{T}$, respectively. In the top graphs of Figure 2, we show that removing the speculated communication channel from the inhibition heads almost always results in a significant decrease in the model’s ability to pass the inhibition signal, with the exception that changing 7.3 does not have a strong effect on its own. In general removing the single component with the highest composition score reduces the Inhibition score by 7-14%, and it is important to consider that this is only when changing a single component in one head at a time. We perform additional experiments with removing/modifying multiple of these components in Section 4 and Appendix D. Thus, we have both behavioral and weight based evidence converging on the interpretation of these subspaces as meaningful communicatoin channels.
@@ 4.1 Interventions on Communication Channels
Figure 4: We find that the 1D inhibition components and 2D duplicate token components finely control which name is avoided by the mover head. On the top, we can selectively inhibit either the first or second name depending on how we scale a vector lying on the 8.10.1 output space. This is strictly controlling relative position. On the bottom, we find that adding or removing duplicate token information from the duplicate channel at the IO or S1 tokens also effectively modulates which name is inhibited. Neither random heads, nor non-communication channel components exhibit these same effects (right). See Appendix D for results on other heads.
@@
Although we get fine control over the attention of the mover head, we have not answered whether this has a real effect on the output behavior. Additionally, Makelov et al. [2024] argue that interventions on subspaces, such as the case here, are prone to a subspace illusion in which the effect is not what it seems. To address this, we measure the Fractional Logit Difference Decrease (FLDD) and Interchange Accuracy of patching the subspaces in the inhibition channel between minimally different IOI examples. GPT2-Small achieves an FLDD of 97.5% and and interchange accuracy of 35%. Such strong results from only patching three 1D subspaces supports the view that these subspaces are a primary mechanism controlling name selection. On OpenWebText [Gokaslan and Cohen, 2019], we find that the inhibition heads are primarily active in lists and settings where repetitions should be avoided, such as “serotonin, dopamine, and…" where the heads will attend from the commas to the previous chemicals. A natural followup given the IOI results and these observations is to explore their role as an indexing function, which we do in the Laundry List task.
5 The Inhibition Channel is Crucial in Token Indexing Tasks
@@ 5.1 Laundry List Task
We propose a synthetic task that is designed to activate the inhibition heads, but allows us to test their effect on an arbitrary number of candidate tokens. An example is given in Figure 1 and more details on how we generated the data are in Appendix E. Each example first mentions N objects, then N-1, and the model must predict the missing item. We create a dataset of 250 prompts for each value of N from 3 to 20.
@@
We find that the inhibition heads are highly active on this task when predicting the last object, like the analysis in Section 4 would predict. The heads attend to and inhibit the occurrences of all of the repeated object tokens.

### v2

@@ Abstract
Although it is known that transformer language models (LMs) pass features from early layers to later layers, it is not well understood how this information is represented and routed by the model. We analyze a mechanism used in two LMs to selectively inhibit items in a context in one task, and find that it underlies a commonly used abstraction across many context-retrieval behaviors. Specifically, we find that models write into low-rank subspaces of the residual stream to represent features which are then read out by later layers, forming low-rank communication channels [Elhage et al., 2021] between layers. A particular 3D subspace in model activations in GPT-2 can be traversed to positionally index items in lists, and we show that this mechanism can explain an otherwise arbitrary-seeming sensitivity of the model to the order of items in the prompt. That is, the model has trouble copying the correct information from context when many items “crowd" this limited space. By decomposing attention heads with the Singular Value Decomposition (SVD), we find that previously described interactions between heads separated by one or more layers can be predicted via analysis of their weight matrices alone. We show that it is possible to manipulate the internal model representations as well as edit model weights based on the mechanism we discover in order to significantly improve performance on our synthetic Laundry List task, which requires recall from a list, often improving task accuracy by over 20%. Our analysis reveals a surprisingly intricate interpretable structure learned from language model pretraining, and helps us understand why sophisticated LMs sometimes fail in simple domains, facilitating future analysis of more complex behaviors. [^1] https://github.com/jmerullo/talking_heads.git
@@ 1 Introduction
Despite the impressive capabilities of LMs, they often suffer from seemingly arbitrary sensitivities to prompts. These failure cases are particularly troubling because they are not systematic; it is very difficult to predict when, for example, the order of information seemingly randomly causes a model to fail [Pezeshkpour and Hruschka, 2023, Liu et al., 2024, Li and Gao, 2024, Zheng et al., 2024, Zhou et al., 2023], or the format of a prompt hurts performance [Liu et al., 2023, Sclar et al., 2023, Zhao et al., 2021, Lu et al., 2022, Webson and Pavlick, 2022]. As LLMs become increasingly ubiquitous, we will require more principled ways of anticipating and remedying unstable or unwanted behaviors [Yu et al., 2024, Yong et al., 2023]. Understanding the mechanisms in play within LLMs, and connecting those mechanisms to behavior, could enable such principled approaches.
@@
One aim of interpretability research is to explain model behaviors, so is it possible to explain why some particular failure exists? In this paper, we consider a simple laundry list task that exhibits one such undesirable instability. Specifically: Transformer language models (LMs) struggle to reliably recall items from a list as the length of the list increases, and performance can vary wildly depending on the position of the item in the list that is being recalled (Figure 1). This instability is not obvious from the model architecture itself–i.e., unlike their predecessors Elman [1991], Transformers [Vaswani et al., 2017] can use attention to recall freely from anywhere in context. Thus, we use this task as a case study in order to connect the low-level emergent mechanisms which are encoded during LM pretraining to observable behavior, and illustrate as a proof of concept that a precise mechanistic understanding of LMs can be used to explicate and, perhaps, remedy model performance in practice.
@@
Specifically, by building on recent work in circuit analysis [Elhage et al., 2021, Wang et al., 2022, Goldowsky-Dill et al., 2023, Quirke and Barez, 2024, Merullo et al., 2024, Hanna et al., 2023], we demonstrate how a Transformer LM (GPT2 small, Pythia 160m) passes information from early layers to later ones using low-rank subspaces. These communication channels are proposed in Elhage et al. [2021], but understanding their implementation in weights is an open challenge [Makelov et al., 2024]. This conflicts with circuit analysis, which supports the interpretation of specialization and communication between specific transformer components. Elhage et al. [2021] introduce a score for measuring how much weight matrices in a transformer communicate, but outside of toy models, it has been difficult to interpret the results [Singh et al., 2024]. We build on this work by proposing a method using the Singular Value Decomposition to find these channels by decomposing one matrix into all of its component signals, and find it is much more interpretable than the non-decomposed variant.We focus on two examples of such channels (inhibition and duplicate detection, §2), and find that they are are very low-rank (1 or 2 dimensions), easily interpretable, and causally important for specific model behaviors. We also show this method can be used to perform model editing at training time (Appendix H.2), and provide some encouraging early evidence we can perform circuit discovery without running models (Appendix I). Specifically, our contributions are as follows:
@@
  • We explore a fundamental question in interpretability on how information passed from layer to layer in an LM is represented internally. We find “communication channels" [Elhage et al., 2021] encoded in the weights that connect attention heads separated by several layers.
  • We propose a simple extension to previous weight-based methods that more effectively isolates low-rank signals passed through such channels.
@@
  • We show that low-rank communication mechanism plays a role in prompt sensitivity on an item recall task that otherwise seems idiosyncratic, and that intervention in the mechanism can be used to affect downstream behavior to improve performance.
@@
Our findings indicate more broadly that LMs are capable of learning intricately structured representations from self-supervised pretraining without inductive biases. This may have important implications for the emergence of abstract and content-independent operations, and for developing methods for steering and understanding these models (see Discussion, §7).
@@
Figure 1: Language models are often sensitive to arbitrary changes in a prompt, for example the order in which objects are listed (right). This problem is more pronounced as the number of objects increases (left) even though it is not obvious where the issue stems from in the model. We broadly explore how information is routed through a model and focus on a mechanism that is in part responsible for this (in)ability.
@@ 2 Background
Throughout our work, we consider decoder-only transformer language models and primarily use GPT2-Small. The modern attention mechanism used by these models Vaswani et al. [2017] uses multiheaded attention, where Query and Keys control which tokens from earlier in context are attended to and Value and Output matrices control what information is moved from these tokens. An important abstraction we use in our work relies on rewriting these matrices as the products QK (Query*Key) and OV (Output*Value) [^2] Following convention, we refer to this matrix as OV, but the Transformer Lens library implements right-hand matrix multiplication so we actually use VO. This does not effect our results. For an individual head, these matrices are themselves low-rank compared to the embedding dimension of the network (due to down-projecting the input e.g., they have size 768x64 dimensions in GPT-2). This is a useful property as we are motivated by looking for subspaces that are written into/read from by these matrices and this reduces the search space.
@@
We build on work from IOI [Wang et al., 2022] which documents a circuit that appears in multiple tasks [Merullo et al., 2024]. This circuit includes mover heads, which copy tokens from context to the output, and inhibition heads which (optionally) block the mover heads’ attention to certain tokens and thus prevent certain tokens from being copied. Inhibition heads are known to receive signals from duplicate token head value vectors which help inform which tokens to inhibit. In GPT2-Small, the known inhibition heads are 7.3, 7.9, 8,6, and 8.10 and we consider their communication to the mover head 9.9333for simplicity we only consider this mover head, but we do not find the choice matters much.. This is the example we use for query composition experiments. In Section 5 we explore this circuit’s role in problems of prompt sensitivity and a learned structure to control the indexing of tokens in the context window.
@@ 3 Identifying Communication Channels in Low-Rank Subspaces
In this section, we test the hypothesis that model components like attention heads communicate through signals in low-rank subspaces of the residual stream and that we can find these signals in the weights themselves. We investigate one case of each type of composition outlined in Section 2 and find positive evidence for this hypothesis with query and value composition in 1 and 2 dimensional subspaces, but not key composition (see Appendix C). Because we are able to localize the query and value signals to such small representation spaces, we find that we are able to interpret and control these features with intervention experiments in Section 4.
@@ 3.1 Composition Score
The Composition Score (CS) introduced in Elhage et al. [2021] is a weight based metric of how much two weight matrices ‘talk’ to each other when they are separated by layers. That is, $\mathbf{W_{1}}$ might write information into a subspace in the residual stream that is read out by $\mathbf{W_{2}}$. In Query composition for example, $\mathbf{W_{1}}$ is the OV matrix of some head, and $\mathbf{W_{2}}$ is the QK matrix.
@@
We take advantage of the fact that circuit analysis in works like Wang et al. [2022] tells us that, for example, head 3.0 (duplicate head) interacts with head 7.9 (inhibition head) through value composition and 7.9 with 9.9 (mover head) in query composition.
@@ 3.2 Composition with Decomposed Component Matrices
We initially use the composition score in Equation 1 to try predict these interactions in the weights, but find these results are largely noisy and uninterpretable. This is briefly demonstrated in Appendix I. We find that despite empirically knowing interactions exist between heads, we do not find they reliably have higher composition scores than any random head. Although the composition score has been shown to be useful on small toy models [Elhage et al., 2021], previous work has also shown that on larger models, the signal the composition score conveys is extremely noisy [Singh et al., 2024]. Therefore, we turn to the Singular Value Decomposition (SVD), defined as $\mathbf{W}=\mathbf{USV^{T}}$, on the QK and OV matrices to decompose the attention heads into orthogonal components which determine the input and output spaces of the matrix. This allows us to individually view subspaces read from/written into ordered by the amount of variance of the transformation of the matrix they account for. This helps us answer our original hypothesis that model components communicate across layers in low-rank subspaces of the residual stream.
@@
We find that decomposing weight matrices this way is very effective at cleaning up the composition score signal. We find attention heads that have very high relative composition scores with one component matrix of another head. For example, the second component of head 8.6 (referred to as 8.6.2) composes far more with mover head 9.9 than any of the other 63 components in 8.6 (5 standard deviations higher than the average) or when considering the full matrix as in $CS(\mathbf{OV}^{8.6},\mathbf{QK}^{9.9})$. The bottom graphs in Figure 2 show these results for the inhibition heads. All of the inhibition head exhibit a similar phenomenon of single component dominance. The duplicate token head 3.0 also value-composes with inhibition heads in a similar way, using two components (3.0.1 and 3.0.2) far more than any other. Results are shown in Appendix D.
We can also use this decomposition to find specific pairs of heads that talk more than others. With the knowledge that two heads talk through a specific component of one head, we can find the other heads that communicate through this pathway. Doing so lets us find the signal encoding almost the full IOI circuit in GPT2-Small directly from the weights, without running the model. We outline these results in Appendix I.
@@ 3.3.2 Results
Our editing technique is simply to zero out one component at a time test how this affects copying behavior across a dataset of IOI examples. One way to think about this is zeroing out one singular value of e.g., the OV or matrix, or subtracting one of the component matrices from the sum in Equation 2. We must make the edit to the decomposition and then split the matrices back out so that we can run the model. Given $\mathbf{OV}=\mathbf{USV}^{T}$, after zeroing out some singular value in $\mathbf{S}$ (forming $\mathbf{S^{\prime}}$) we can set the Output and Value matrices to be $\mathbf{U}\sqrt{\mathbf{S^{\prime}}}$ and $\sqrt{\mathbf{S^{\prime}}}\mathbf{V}^{T}$, respectively. In the top graphs of Figure 2, we show that removing the speculated communication channel from the inhibition heads almost always results in a significant decrease in the model’s ability to pass the inhibition signal, with the exception that changing 7.3 does not have a strong effect on its own. In general removing the single component with the highest composition score reduces the Inhibition score by 7-14%, and it is important to consider that this is only when changing a single component in one head at a time. We perform additional experiments with removing/modifying multiple of these components in Section 4 and Appendix E. Thus, we have both behavioral and weight based evidence converging on the interpretation of these subspaces as meaningful communication channels.
@@ 4.1 Interventions on Communication Channels
Figure 4: We find that the 1D inhibition components and 2D duplicate token components finely control which name is avoided by the mover head. On the top, we can selectively inhibit either the first or second name depending on how we scale a vector lying on the 8.10.1 output space. This is strictly controlling relative position. On the bottom, we find that adding or removing duplicate token information from the duplicate channel at the IO or S1 tokens also effectively modulates which name is inhibited. Neither random heads, nor non-communication channel components exhibit these same effects (right). See Appendix E for results on other heads.
@@
Although we get fine control over the attention of the mover head, we have not answered whether this has a real effect on the output behavior. Additionally, Makelov et al. [2024] argue that interventions on subspaces, such as the case here, are prone to a subspace illusion in which the effect is not what it seems. To address this, we measure the Fractional Logit Difference Decrease (FLDD) and Interchange Accuracy of patching the subspaces in the inhibition channel between minimally different IOI examples. GPT2-Small achieves an FLDD of 97.5% and and interchange accuracy of 35%. Makelov et al. [2024] report a baseline that achieves -8% FLDD and 0.0% Interchange. By taking the gradient to directly optimize a single vector for this task they achieve 111.5% FLDD and 45.1% Interchange. Our results in comparison, support the view that these 1D subspaces are a primary mechanism controlling name selection. In Appendix G we also find that these inhibition signals are active broadly during next work prediction. On OpenWebText [Gokaslan and Cohen, 2019], we find that the inhibition heads are primarily active in lists and settings where repetitions should be avoided; for example, in comma separated lists (attending from commas to previously seen items). A natural followup given the IOI results and these observations is to explore their role as an indexing function, which we perform with the Laundry List task.
5 The Inhibition Channel is Crucial in Context Retrieval Tasks
@@ 5.1 Laundry List Task
We propose a synthetic task that is designed to activate the inhibition heads, but allows us to test their effect on an arbitrary number of candidate tokens. An example is given in Figure 1 and more details on how we generated the data are in Appendix F. Each example first mentions N objects, then N-1, and the model must predict the missing item. We create a dataset of 250 prompts for each value of N from 3 to 20.
@@
We find that the inhibition heads are highly active (large attention on non EOS tokens) on this task when predicting the last object, like the analysis in Section 4 would predict. The heads attend to and inhibit the occurrences of all of the repeated object tokens.